Rewrite the expression log5(√(x^3+1)^3(x^2+5)^4/(x^2+4x+2)^5))^3


1
Expert's answer
2022-06-13T14:32:34-0400
"\log_5(\dfrac{\sqrt{(x^3+1)^3(x^2+5)^4}}{(x^2+4x+2)^5})^3"

"=\dfrac{9}{2}\log_5(x^3+1)+6\log_5(x^2+5)"

"-15\log_5(x^2+4x+2))"